LUCIFER Training & Preparations for Partner Science:
UT 2010 July 02

Observer: DThompson, KRueff, PHibon
Telescope Operator: SAllanson
Instrument Support: TShih
Telescope Support: JUrban
SW support: JKraus (on call)

Summary

Details

04:27 Afternoon unsettled weather, delaying opening until now (MST 9:30). Tout=12.8 DewPt=7.1 RH=68.6% Wind ~5.5m/s Tmirror1=12.2, mostly clear. Continuing with GCS binning set to 2x2 by default.

Correct initial collimation/pointing:
  • 04:37 Preset to ACT0211 to correct pointing
  • 04:43 Preset to BS9160 to collimate, off-axis GS

INAF_Grazian MOS mask LST=16:11 (same as yesterday) once done with collimating, too late to take a full set.

04:50 Preset to WT10_356 to correct pointing before spectroscopic acquisition. luci#0005

Thompson: HR10 1645+46
  • Note: Pascale's suggestion, zoom in on skycat (2x is good) before auto-set cut levels
  • Acq at 04:53, Sky #0006, Src#0007 slit#0008 through slit#0009...
  • 05:08 HR10_Hspec (10x300s MER), start at EL~76, FWHM on guider ~1.0 arcsec and fluctuating Files 0012-0021
  • 06:04 insert calib unit.
  • 06:05 specflatH files off=0022 on=0023 off=0024 on=0025
  • 06:08 specarcH Xe: off=26 on=0027, Ar: on=28 off=29
  • 06:12 34Her_acq
  • 06:19 34Her_Hspec files 0033, 0034

Kuhn: HS1946+7658
  • 06:23 Executing HIP102208_HKspec at 34Her, source in the L450 slit. Files 0035, 0036
  • 06:32 HS1946KH_acq 0037=sky, 0038=src, 0039=slit 0040,0041=throughslit Note: clouds approaching, RH rising!
  • 06:43 HS1946HK_sci files: 0042-0045, guider showing ~0.85 arcsec FWHM
  • 07:02 lost guide star about 2 minutes into last exposure.
  • 07:06 closing dome, last spectrum significantly weaker than the other three.
  • 07:13 specflatHK in closed dome, files off=46-48, on=49-51
  • 07:15 specarcHK in closed dome, files Xe: off=52 on=53, Ar off=54 on=55

Same place, check some filters spectroscopically (LS450 slit):
  • FeII.specflat, files off=56 on=57
  • H2.specflat, files off=58 on=59
  • BrG.specflat, files off=60 on=61
  • Ks.specflat, files off=62 on=63

08:15 RH94% with rising dewpoint and falling temps. Sigh.

10:50 Re-opening for the last hour
  • 10:58 Preset to ACT0398 to correct pointing near Olga's telluric. #0064
  • 11:01 HIP102208_acq #0065=src #0066=slit 0067, 68 through slit
  • 11:13 HIP102208_HKspec files #0069, 0070 (I hope this works)

Pedani: 21h 0d
  • 11:17 Preset to 2120+0006, Ks 900s files: 0071-0085
  • Already see background on wfs. Seeing ~2", sorry Marco.

11:30 Guide star lost, lucifer image shows systematic jumps. Closing.

Co-pointing Model observations (28 points) NOT ABLE TO TAKE THESE DATA!
  • Change to IRTC authorized on right size (telescope and LBTtools)
  • Correct pointing and initial collimation
  • Start pointing log
  • Preset with IRTC, GUIDE mode gs=0 at PA=0
  • Take images with IRTC and LUCIFER to record source, timestamp

Morning Darks Start with file 0076

-- DavidThompson - 02 Jul 2010
Topic revision: r2 - 02 Jul 2010, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback